OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 A 8 – 4x B 8 + 4x C –8 + 4x D –8 − 4x

OpenStudy (anonymous):

Solve for y in the first equation y=4x-8

OpenStudy (texaschic101):

4x - y = 8 -y = 8 - 4x y = 4x - 8

OpenStudy (anonymous):

So it would be A

OpenStudy (anonymous):

More like C

OpenStudy (texaschic101):

no...it is -8 + 4x

OpenStudy (anonymous):

ohhhhhh okay i understand now :)

OpenStudy (texaschic101):

that is good that you understand :)

OpenStudy (anonymous):

thanks for all your help :)

OpenStudy (texaschic101):

I was not the only one that helped....but your welcome for the part I did :)

OpenStudy (anonymous):

@incomplte thanks for your help ;)

OpenStudy (anonymous):

You guys are so nice lol. We are just here to help

OpenStudy (anonymous):

could you guys help me with some more?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 2x − 4y = 6 3x − y = 12 A Add the two equations together B Multiply the first equation by –4 C Multiply the second equation by –4 D Multiply the second equation by 4

OpenStudy (texaschic101):

multiply 2nd equation by -4.....C

OpenStudy (anonymous):

thanks @texaschic101 :)

OpenStudy (anonymous):

@texaschic101 could you help me with just 2 more?

OpenStudy (texaschic101):

lets have them....

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? A 330 B 510 C 520 D 660

OpenStudy (texaschic101):

a stands for one step equations b stands for two step equations a + b = 840 ---> a = 840 - b 3a - b = 870 now sub 840 - b in for a in the second equation..... do u know how to do this with the information I have given u ?

OpenStudy (anonymous):

to get the answer i would have to subtract 840-870? is that right

OpenStudy (anonymous):

wait i got 510 is that correcT?

OpenStudy (texaschic101):

no.....let me show you.... I set the 2nd equation up wrong.... It is supposed to be 3a - 2b = 870 (sub 840 - b in for a) 3(840 - b) - 2b = 870 2520 - 3b - 2b = 870 2520 - 5b = 870 2520 - 870 = 5b 1650 = 5b 1650/5 = 5b/5 330 = b a + b = 840 a + 330 = 840 a = 840 - 330 a = 510 check... 3a - 2b = 870 3(510) - 2(330) = 870 1530 - 660 = 870 870 = 870 (correct) yes you have it right......sorry I told you wrong about the equation....you must have found my mistake to get the correct answer....good for you.